Packaging Solutions Corporation manufactures and sells a wide variety of packaging products. Performance reports are prepared monthly for each department. The planning budget and flexible budget for the Production Department are based on the following formulas, where q is the number of labor-hours worked in a month:

Cost Formulas
Direct labor $16.40q
Indirect labor $4,100 + $1.50q
Utilities $5,200 + $0.60q
Supplies $1,600 + $0.30q
Equipment depreciation $18,200 + $2.50q
Factory rent $8,200
Property taxes $2,500
Factory administration $13,500 + $0.60q

The Production Department planned to work 4,300 labor-hours in March; however, it actually worked 4,100 labor-hours during the month. Its actual costs incurred in March are listed below:

Actual Cost Incurred in March
Direct labor $ 68,840
Indirect labor $ 9,730
Utilities $ 8,170
Supplies $ 3,100
Equipment depreciation $ 28,450
Factory rent $ 8,600
Property taxes $ 2,500
Factory administration $ 15,310

Required:

1. Prepare the Production Department’s planning budget for the month.

2. Prepare the Production Department’s flexible budget for the month.

3. Calculate the spending variances for all expense items.
Answer #1
1
Planning Budget
Budgeted labor hours 4300
Direct labor 70520 =4200*16.5
Indirect labor 10550 =4100+(4300*1.5)
Utilities 7780 =5200+(4300*0.6)
Supplies 2890 =1600+(4300*0.3)
Equipment Depreciation 28950 =18200+(4300*2.5)
Factory rent 8200
Property taxes 2500
Factory administration 16080 =13500+(4300*0.6)
Total expenses 147470
2
Flexible Budget
Actual labor hours 4100
Direct labor 67240 =4100*16.4
Indirect labor 10250 =4100+(4100*1.5)
Utilities 7660 =5200+(4100*0.6)
Supplies 2830 =1600+(4100*0.3)
Equipment Depreciation 28450 =18200+(4100*2.5)
Factory rent 8200
Property taxes 2500
Factory administration 15960 =13500+(4100*0.6)
Total expenses 143090
3
Actual Spending variance Flexible budget Activity variance Planning budget
Labor hours 4100 4100 4300
Direct labor 68840 1600 U 67240 3280 F 70520
Indirect labor 9730 520 F 10250 300 F 10550
Utilities 8170 510 U 7660 120 F 7780
Supplies 3100 270 U 2830 60 F 2890
Equipment Depreciation 28450 0 None 28450 500 F 28950
Factory rent 8600 400 U 8200 0 None 8200
Property taxes 2500 0 None 2500 0 None 2500
Factory administration 15310 650 F 15960 120 F 16080
Total expenses 144700 1610 U 143090 4380 F 147470